OpenStudy (anonymous):

can somebody help me?????Solve the system. y = -1/3 x + 2 and x + 3y = 3 (0, 1) (1, 0) (3, 1/3) (3/2, 0) no solution

OpenStudy (calculusfunctions):

Substitute equation 1 into equation 2. In other words, replace the y in the second equation with the right side of the first equation. So you have\[x +3(-\frac{ 1 }{ 3 }x +2)=3\]Now solve for x and then substitute the x value back into either equation and solve for y.

OpenStudy (calculusfunctions):

You will find that there is no such x value here.

OpenStudy (anonymous):

wait so you mixed the equations?

OpenStudy (calculusfunctions):

If by mix you mean substitute, then yes.

OpenStudy (anonymous):

so substirtute the possible answers for the equations?

OpenStudy (aripotta):

he substituted for y and then tried to solve for x. however, there is no answer for x. so there's no solution

OpenStudy (calculusfunctions):

Since y is isolated in equation 1, it makes sense to replace the y in equation 2 with whatever y equals in equation 1.

OpenStudy (anonymous):

so you changed the equation to y=mx+b and then solved for x?

OpenStudy (aripotta):

no. in the first equation, y was already solved for in terms of x. so he substituted that into the second equation

OpenStudy (anonymous):

i must sound really stupid but can you like show me the steps that you did that?

OpenStudy (calculusfunctions):

I did above. Did you see?

OpenStudy (aripotta):

x + 3y = 3 x + 3(-1/3 x + 2) = 3

OpenStudy (calculusfunctions):

That's exactly what I did.

OpenStudy (aripotta):

once you distribute... x + 3(-1/3 x + 2) = 3 x - x + 6 = 3 6 = 3 which isn't correct...so there's no answer for x

OpenStudy (calculusfunctions):

Yes, Very good!

OpenStudy (aripotta):

does that make sense, emma?

OpenStudy (anonymous):

i think so thanks peoples im havng trouble figuring out who to give the medal too though

OpenStudy (aripotta):

doesn't matter to me

OpenStudy (calculusfunctions):

x - x + 6 = 3 x - x = 3 - 6 0x = -3 x = -3/0 But a number divided by zero is undefined so there is no solution.

OpenStudy (calculusfunctions):

A medal is not what's important. What matters is do you understand what we just taught you?

OpenStudy (anonymous):

yes much better now:)

OpenStudy (anonymous):

thank you

OpenStudy (calculusfunctions):

Great! Then I feel a sense of accomplishment!

OpenStudy (anonymous):

:)

OpenStudy (calculusfunctions):

Do you need help with anything else?

OpenStudy (calculusfunctions):

I guess not. Have a great holiday and stay safe.

OpenStudy (anonymous):

oh thanks